# Leave-One-Out Analysis for Nonconvex Robust Matrix Completion with General Thresholding Functions

### v1

@@ 5 Induction Steps
In the induction steps, our proofs and the ones in [11] differ mainly in the following three aspects:
  • We use matrix Bernstein’s inequality to bound $\left\|e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{t,m}-L^{\star}\right)V^{t+1,m}\right\|_{2}$, and use Lemma 11 to bound $\tilde{T}_{2a},~{}\tilde{T}_{2b}$, and $\tilde{B}_{1},~{}\tilde{B}_{2}$, resulting in better bounds.
  • We use Lemma 12 to derive the bounds for $\tilde{T}_{2c}$ and $\tilde{B}_{3}$, which reduces the order of $r$ from 5 in [11] to 2.
  • The outlier term $\left\|\mathcal{P}_{\Omega}^{(-m)}\left(S^{t,m}-S^{t}\right)V^{t+1,m}\right\|_{\mathrm{F}}$ needs to be bounded carefully.
@@
Step 1: The Bound For $\left\|E^{t,\infty}\right\|_{2}$. By Lemma 1 with the bounds (3a) and (3b),
@@
$\leq$ $\frac{5C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}\cdot\left(2+\frac{5C_{10}}{C_{0}}\right)\sqrt{\frac{\mu r}{n}}\cdot\left(1+\frac{1}{C_{0}}\right)\sigma_{1}^{\star}+\frac{\mu r}{n}\cdot\frac{7}{C_{0}}\left(\frac{1}{\kappa\mu r}\sigma_{r}^{\star}\right)\gamma^{t}$
$\leq$ $\frac{12C_{10}}{C_{0}}\cdot\frac{1}{\kappa\mu r}\left(\frac{\mu r}{n}\sigma_{1}^{\star}\right)\gamma^{t}\leq\left(\frac{\mu r}{n}\sigma_{1}^{\star}\right)\gamma^{t}.$
@@
$\left\|E_{1}^{t,i}\right\|_{2}=\left\|p^{-1}\mathcal{P}_{\Omega}^{(-i)}\left(S^{t,i}-S^{\star}\right)\right\|_{2}\leq p^{-1}\cdot\left(2\alpha pn\right)\cdot C_{\text{thresh}}\left(\frac{\mu r}{n}\sigma_{1}^{\star}\right)\gamma^{t}\leq\frac{1}{2C_{0}}\cdot\left(\frac{1}{\kappa^{2}\mu r}\sigma_{r}^{\star}\right)\gamma^{t+1}.$
@@
$\leq$ $c_{5}\sqrt{\frac{2rn\log n}{p}}\cdot\frac{12C_{10}}{C_{0}}\frac{\sigma_{r}^{\star}}{n}\gamma^{t}\leq\frac{1}{2C_{0}}\cdot\left(\frac{1}{\kappa^{2}\mu r}\sigma_{r}^{\star}\right)\gamma^{t+1},$
@@
$\left\|E^{t,i}\right\|_{2}\leq\frac{1}{C_{0}}\cdot\left(\frac{1}{\kappa^{2}\mu r}\sigma_{r}^{\star}\right)\gamma^{t+1}.$
@@
$\left\|T_{1}\right\|_{2}+\left\|T_{2}\right\|_{2}+\frac{1}{2}\left\|T_{3}\right\|_{2}\leq\frac{16}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}.$ (13)
@@
$\left\|e_{m}^{T}E_{1}^{t,i}V^{t+1,i}\right\|_{2}\leq\frac{\sigma_{r}^{\star}}{2C_{0}}\cdot\left(\frac{1}{\kappa^{2}\mu r}\gamma^{t+1}\right)\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right),$
bounding $e_{m}^{T}E_{2}^{t,i}V^{t+1,i}$ is technically quite involved. For ease of notation, let $Q:=\left(G^{t+1,m}\right)^{T}G^{t+1,i}$ in this part. We have
$e_{m}^{T}E_{2}^{t,i}V^{t+1,i}=$ $e_{m}^{T}E_{2}^{t,i}V^{t+1,m}Q+e_{m}^{T}E_{2}^{t,i}\left(V^{t+1,i}-V^{t+1,m}Q\right)$
$=$ $e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{t,i}-L^{\star}\right)V^{t+1,m}Q+e_{m}^{T}E_{2}^{t,i}\left(V^{t+1,i}-V^{t+1,m}Q\right)$
$=$ $\underbrace{e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{t,m}-L^{\star}\right)V^{t+1,m}Q}_{\tilde{T}_{1}}+\underbrace{e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{t,i}-L^{t,m}\right)V^{t+1,m}Q}_{\tilde{T}_{2}}$
$+\underbrace{e_{m}^{T}E_{2}^{t,i}\left(V^{t+1,i}-V^{t+1,m}Q\right)}_{\tilde{T}_{3}}.$
@@
$\left\|e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{t,m}-L^{\star}\right)V^{t+1,m}Q\right\|_{2}=\left\|e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{t,m}-L^{\star}\right)V^{t+1,m}\right\|_{2}.$
@@
$\leq$ $C_{10}\cdot\left(\sqrt{\frac{n\log n}{p}}+\frac{\log n}{p}\right)\cdot\frac{12C_{10}}{C_{0}}\frac{\sigma_{r}^{\star}}{n}\gamma^{t}\cdot\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)$
$\leq$ $\frac{C_{10}}{C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\gamma^{t+1}\right)\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right),$
@@
$\leq$ $\frac{\gamma}{9\kappa}\cdot\sqrt{\frac{n}{\mu r}}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right),$
@@
$p\geq\frac{400C_{10}^{2}}{\gamma^{2}}\cdot\frac{\kappa^{2}\mu r\log n}{n}.$
@@
$\left\|\tilde{T}_{2a}\right\|_{2}\leq$ $\frac{8C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}\cdot\left(1+\frac{1}{C_{0}}\right)\sigma_{1}^{\star}\cdot\frac{\gamma}{9\kappa}\sqrt{\frac{n}{\mu r}}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)$
$\leq$ $\frac{C_{10}}{C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\gamma^{t+1}\right)\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right).$
@@
$\left\|S^{t,m,i}\right\|_{2}=\left\|\Sigma^{t,m}G^{t,i,m}-G^{t,i,m}\Sigma^{t,i}\right\|_{2}\leq 4\kappa\cdot\left\|E^{t-1,i}-E^{t-1,m}\right\|_{2}\leq\frac{8}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sigma_{r}^{\star}\right)\gamma^{t}.$
@@
$\tilde{T}_{2b}\leq\frac{2}{C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa^{2}\mu r}\gamma^{t+1}\right)\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right).$
@@
$\leq$ $\frac{1}{C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\gamma^{t+1}\right)\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right),$
@@
$p\geq\frac{100c_{5}^{2}C_{10}^{2}}{\gamma^{2}}\cdot\frac{\kappa^{2}\mu^{2}r^{2}\log n}{n}.$
$\bullet$ Bounding $\tilde{T}_{3}$. Recalling that $Q=\left(G^{t+1,m}\right)^{T}G^{t+1,i}$,
$\left\|\tilde{T}_{3}\right\|_{2}=$ $\left\|e_{m}^{T}E_{2}^{t,i}\left(V^{t+1,i}-V^{t+1,m}\left(G^{t+1,m}\right)^{T}G^{t+1,i}\right)\right\|_{2}$
$\leq$ $\left\|e_{m}^{T}E_{2}^{t,i}\left(V^{t+1,i}-V^{\star}G^{t+1,i}\right)\right\|_{2}+\left\|e_{m}^{T}E_{2}^{t,i}\left(V^{\star}-V^{t+1,m}\left(G^{t+1,m}\right)^{T}\right)G^{t+1,i}\right\|_{2}$
$=$ $\left\|e_{m}^{T}E_{2}^{t,i}\Delta_{V}^{t+1,i}\right\|_{2}+\left\|e_{m}^{T}E_{2}^{t,i}\Delta_{V}^{t+1,m}\right\|_{2}.$
The two terms can be bounded similarly. For example, on the event that $\sum_{k=1}^{n}\delta_{mk}\leq 2pn,$ a direct calculation yields that
$\left\|e_{m}^{T}E_{2}^{t,i}\Delta_{V}^{t+1,i}\right\|_{2}=$ $\left\|e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{t,i}-L^{\star}\right)\Delta_{V}^{t+1,i}\right\|_{2}$
$\leq$ $\sum_{k=1}^{n}\left|1-\delta_{mk}/p\right|\cdot\left\|L^{t,i}-L^{\star}\right\|_{\infty}\cdot\left\|\left(\Delta_{V}^{t+1,i}\right)_{k,:}\right\|_{2}$
$\leq$ $3n\cdot\left\|L^{t,i}-L^{\star}\right\|_{\infty}\cdot\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}$
$\leq$ $3n\cdot\frac{12C_{10}}{C_{0}}\frac{\sigma_{r}^{\star}}{n}\gamma^{t}\cdot\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\leq\frac{36C_{10}}{C_{0}}\sigma_{r}^{\star}\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}.$
@@
$\left\|e_{m}^{T}E_{2}^{t,i}V^{t+1,i}\right\|_{2}\leq$ $\frac{3+2C_{10}}{C_{0}}\left(\frac{\sigma_{r}^{\star}}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}+\frac{3+74C_{10}}{C_{0}}\sigma_{r}^{\star}\left\|\Delta^{t+1,\infty}\right\|_{2,\infty},$
@@
$\left\|T_{4}\right\|_{2}\leq$ $\frac{4+3C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}+\frac{4+80C_{10}}{C_{0}}\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}.$
@@
$\left(1-\frac{4+80C_{10}}{C_{0}}\right)\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\leq$ $\frac{20+3C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}.$
When $C_{10}\geq 20$, and $C_{0}\geq 5(4+80C_{10})$, it reduces to
$\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\leq\frac{5C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}.$
@@
$\sum_{k=1}^{n}w_{ik}\left(V^{t+1,m}_{kj}\right)^{2}=\left\|\mathcal{P}_{\Omega_{S^{\star}}^{(-m)}}\left(e_{i}e_{j}^{T}\cdot\left(V^{t+1,m}\right)^{T}\right)\right\|_{\mathrm{F}}^{2}\leq\left\|\mathcal{P}_{\Omega_{S^{\star}}}\left(e_{i}e_{j}^{T}\cdot\left(V^{t+1,m}\right)^{T}\right)\right\|_{\mathrm{F}}^{2},$
and we will show that Lemma 13 can be applied to bound this term. One problem is that when $\alpha$ is very small, $\Omega_{S^{\star}}$ may be too sparse for Lemma 13 to work. Thus as inspired by [9], we consider an equivalent model that generates $\Omega_{S^{\star}}$: (1) sample $\Omega$ from the i.i.d. Bernoulli model with parameter $p$; (2) generate $\Omega^{\prime}\subseteq\Omega$ such that each $(i,j)\in\Omega$ is included in $\Omega^{\prime}$ independently with probability $\alpha^{\prime}~{}(\alpha^{\prime}\geq\alpha)$; (3) for each $(i,j)\in\Omega^{\prime}$, it is included in $\Omega_{S^{\star}}$ independently with probability $\frac{\alpha}{\alpha^{\prime}}$. As a result, $\Omega_{S^{\star}}\subseteq\Omega^{\prime}$, and we get the freedom to choose a proper value for $\alpha^{\prime}$. Note that
$\left\|U^{t+1,m}\left(G^{t+1,m}\right)^{T}\left(\Sigma^{t+1,m}\right)^{1/2}-U^{\star}\left(\Sigma^{\star}\right)^{1/2}\right\|_{2,\infty}$ (18)
$\leq$ $\left\|\left(U^{t+1,m}\left(G^{t+1,m}\right)^{T}-U^{\star}\right)\left(\Sigma^{t+1,m}\right)^{1/2}\right\|_{2,\infty}+\left\|U^{\star}\left(\left(\Sigma^{t+1,m}\right)^{1/2}-\left(\Sigma^{\star}\right)^{1/2}\right)\right\|_{2,\infty}$
$\leq$ $\left\|\Delta^{t+1,m}\right\|_{2,\infty}\cdot\left(\left(1+\frac{1}{C_{0}}\right)\sigma_{1}^{\star}\right)^{1/2}+\sqrt{\frac{\mu r}{n}}\cdot\frac{\left\|E^{t,m}\right\|_{2}}{\left(\left(1-\frac{1}{C_{0}}\right)\sigma_{r}^{\star}\right)^{1/2}+\left(\sigma_{r}^{\star}\right)^{1/2}}$
$\leq$ $\frac{5C_{10}}{C_{0}}\sqrt{\left(1+\frac{1}{C_{0}}\right)}\cdot\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}\cdot\left(\sigma_{1}^{\star}\right)^{1/2}+\frac{1}{C_{0}}\cdot\left(\frac{1}{\kappa^{2}\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}\cdot\left(\sigma_{r}^{\star}\right)^{1/2}$
$\leq$ $\frac{6C_{10}}{C_{0}}\cdot\frac{1}{\kappa\sqrt{n}}\left(\sigma_{1}^{\star}\right)^{1/2}\leq\frac{c_{13}}{\kappa\sqrt{n}}\left(\sigma_{1}^{\star}\right)^{1/2},$
where $c_{13}>0$ is the small universal constant from Lemma 13, and the last inequality holds provided that $C_{0}\geq\frac{6C_{10}}{c_{13}}$. Since $\left\|V^{t+1,m}\left(G^{t+1,m}\right)^{T}\left(\Sigma^{t+1,m}\right)^{1/2}-V^{\star}\left(\Sigma^{\star}\right)^{1/2}\right\|_{2,\infty}$ can be bounded in the same way, and $e_{i}e_{j}^{T}\cdot\left(V^{t+1,m}\right)^{T}$ is in the tangent space at
$U^{t+1,m}\left(G^{t+1,m}\right)^{T}\cdot\Sigma^{t+1,m}\cdot\left(V^{t+1,m}\left(G^{t+1,m}\right)^{T}\right)^{T},$
by Lemma 13,
$\left\|\mathcal{P}_{\Omega_{S^{\star}}}\left(e_{i}e_{j}^{T}\cdot\left(V^{t+1,m}\right)^{T}\right)\right\|_{\mathrm{F}}^{2}\leq\left\|\mathcal{P}_{\Omega^{\prime}}\left(e_{i}e_{j}^{T}\cdot\left(V^{t+1,m}\right)^{T}\right)\right\|_{\mathrm{F}}^{2}\leq 40(\alpha^{\prime}p)\kappa\cdot\left\|V^{t+1,m}_{:,j}\right\|_{2}^{2}=40\alpha^{\prime}p\kappa,$
where the second inequality holds as long as $\alpha^{\prime}\geq C_{13}\frac{\mu r\log n}{np}$. Therefore,
@@
$\leq$ $\sum_{i=1}^{n}\left[\sum_{k=1}^{n}w_{ik}\left(S^{t,m}-S^{t}\right)_{ik}^{2}\right]\cdot\left(40\alpha^{\prime}p\kappa r\right)$
$=$ $40\alpha^{\prime}p\kappa r\cdot\left\|\mathcal{P}_{\Omega_{S^{\star}}}^{(-m)}\left(S^{t,m}-S^{t}\right)\right\|_{\mathrm{F}}^{2}.$

### v2

@@ 5 Induction Steps
In the induction steps, our proofs and the ones in [11] differ mainly in the following two aspects:
  • Instead of first bounding $\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}$ and then bounding $\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}$ based on $\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}$, we have established two inequalities (17) and (21) which involve $\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}$ and $\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}$ simultaneously. This allows us to relax the requirement on the distance to the ground truth. Through further refined analysis, in particular by using Lemma 11 to bound $\tilde{T}_{2a},~{}\tilde{T}_{2b}$, and $\tilde{B}_{1},~{}\tilde{B}_{2}$, and using Lemma 12 to derive the bounds for $\tilde{T}_{2c}$ and $\tilde{B}_{3}$, we have improved the sample complexity substantially.
  • The outlier term $\left\|\mathcal{P}_{\Omega}^{(-m)}\left(S^{t,m}-S^{t}\right)V^{t+1,m}\right\|_{\mathrm{F}}$ needs to be bounded carefully, with the help of the Lipschitz property of the thresholding function and Lemma 13.
@@
Step 1: The Bound For $\left\|E^{t,\infty}\right\|_{2}$. Using Lemma 1 with the bounds (3a) and (3b),
@@
$\leq$ $\frac{5C_{10}}{C_{0}}\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}\cdot\left(2+\frac{5C_{10}}{C_{0}}\right)\sqrt{\frac{\mu r}{n}}\cdot\left(1+\frac{1}{C_{0}}\right)\sigma_{1}^{\star}+\frac{\mu r}{n}\cdot\frac{7}{C_{0}}\left(\frac{1}{\mu r}\sigma_{r}^{\star}\right)\gamma^{t}$
$\leq$ $\frac{12C_{10}}{C_{0}}\cdot\frac{1}{\mu r}\left(\frac{\mu r}{n}\sigma_{1}^{\star}\right)\gamma^{t}\leq\left(\frac{\mu r}{n}\sigma_{1}^{\star}\right)\gamma^{t}.$
@@
$\left\|E_{1}^{t,i}\right\|_{2}=\left\|p^{-1}\mathcal{P}_{\Omega}^{(-i)}\left(S^{t,i}-S^{\star}\right)\right\|_{2}\leq p^{-1}\cdot\left(2\alpha pn\right)\cdot C_{\text{thresh}}\left(\frac{\mu r}{n}\sigma_{1}^{\star}\right)\gamma^{t}\leq\frac{1}{2C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sigma_{r}^{\star}\right)\gamma^{t+1}.$
@@
$\leq$ $c_{5}\sqrt{\frac{2rn\log n}{p}}\cdot\frac{12C_{10}}{C_{0}}\frac{\sigma_{1}^{\star}}{n}\gamma^{t}\leq\frac{1}{2C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sigma_{r}^{\star}\right)\gamma^{t+1},$
@@
$\left\|E^{t,i}\right\|_{2}\leq\frac{1}{C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\sigma_{r}^{\star}\right)\gamma^{t+1}.$
@@
$\left\|T_{1}\right\|_{2}+\left\|T_{2}\right\|_{2}+\frac{1}{2}\left\|T_{3}\right\|_{2}\leq\frac{16}{C_{0}}\cdot\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}.$ (13)
@@
$\left\|e_{m}^{T}E_{1}^{t,i}V^{t+1,i}\right\|_{2}\leq\frac{\sigma_{r}^{\star}}{2C_{0}}\cdot\left(\frac{1}{\kappa\mu r}\gamma^{t+1}\right)\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right),$
bounding $e_{m}^{T}E_{2}^{t,i}V^{t+1,i}$ is technically quite involved. First, one has
$e_{m}^{T}E_{2}^{t,i}V^{t+1,i}=$ $e_{m}^{T}E_{2}^{t,i}V^{t+1,m}G^{t+1,i,m}+e_{m}^{T}E_{2}^{t,i}\left(V^{t+1,i}-V^{t+1,m}G^{t+1,i,m}\right)$
$=$ $e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{t,i}-L^{\star}\right)V^{t+1,m}G^{t+1,i,m}+e_{m}^{T}E_{2}^{t,i}\left(V^{t+1,i}-V^{t+1,m}G^{t+1,i,m}\right)$
$=$ $\underbrace{e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{t,m}-L^{\star}\right)V^{t+1,m}G^{t+1,i,m}}_{\tilde{T}_{1}}+\underbrace{e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{t,i}-L^{t,m}\right)V^{t+1,m}G^{t+1,i,m}}_{\tilde{T}_{2}}$
$+\underbrace{e_{m}^{T}E_{2}^{t,i}\left(V^{t+1,i}-V^{t+1,m}G^{t+1,i,m}\right)}_{\tilde{T}_{3}}.$
@@
$\left\|e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{t,m}-L^{\star}\right)V^{t+1,m}G^{t+1,i,m}\right\|_{2}=\left\|e_{m}^{T}\mathcal{H}_{\Omega}^{(-i)}\left(L^{t,m}-L^{\star}\right)V^{t+1,m}\right\|_{2}.$
@@
$\leq$ $C_{10}\cdot\left(\sqrt{\frac{n\log n}{p}}+\frac{\log n}{p}\right)\cdot\frac{12C_{10}}{C_{0}}\frac{\sigma_{1}^{\star}}{n}\gamma^{t}\cdot\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)$
$\leq$ $\frac{C_{10}}{C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\mu r}\gamma^{t+1}\right)\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right),$
@@
$\leq$ $\frac{\gamma}{9}\cdot\sqrt{\frac{n}{\mu r}}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right),$
@@
$p\geq\frac{400C_{10}^{2}}{\gamma^{2}}\cdot\frac{\mu r\log n}{n}.$
@@
$\left\|\tilde{T}_{2a}\right\|_{2}\leq$ $\frac{8C_{10}}{C_{0}}\left(\frac{1}{\kappa\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t}\cdot\left(1+\frac{1}{C_{0}}\right)\sigma_{1}^{\star}\cdot\frac{\gamma}{9}\sqrt{\frac{n}{\mu r}}\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right)$
$\leq$ $\frac{C_{10}}{C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\mu r}\gamma^{t+1}\right)\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right).$
@@
$\left\|S^{t,m,i}\right\|_{2}=\left\|\Sigma^{t,m}G^{t,i,m}-G^{t,i,m}\Sigma^{t,i}\right\|_{2}\leq 4\kappa\cdot\left\|E^{t-1,i}-E^{t-1,m}\right\|_{2}\leq\frac{8}{C_{0}}\cdot\left(\frac{1}{\mu r}\sigma_{r}^{\star}\right)\gamma^{t}.$
@@
$\tilde{T}_{2b}\leq\frac{2}{C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\mu r}\gamma^{t+1}\right)\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right).$
@@
$\leq$ $\frac{1}{C_{0}}\cdot\left(\frac{\sigma_{r}^{\star}}{\mu r}\gamma^{t+1}\right)\left(\sqrt{\frac{\mu r}{n}}+\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\right),$
@@
$p\geq\frac{100c_{5}^{2}C_{10}^{2}}{\gamma^{2}}\cdot\frac{\mu^{2}r^{2}\log n}{n}.$
$\bullet$ Bounding $\tilde{T}_{3}$.
$\left\|\tilde{T}_{3}\right\|_{2}=$ $\left\|e_{m}^{T}E_{2}^{t,i}\left(V^{t+1,i}-V^{t+1,m}G^{t+1,i,m}\right)\right\|_{2}\leq\left\|E_{2}^{t,i}\right\|_{2}\left\|D_{V}^{t+1,i,m}\right\|_{\mathrm{F}}\leq\frac{1}{2C_{0}}\sigma_{r}^{\star}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}.$
@@
$\left\|e_{m}^{T}E_{2}^{t,i}V^{t+1,i}\right\|_{2}\leq$ $\frac{3+2C_{10}}{C_{0}}\left(\frac{\sigma_{r}^{\star}}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}+\frac{3+2C_{10}}{C_{0}}\sigma_{r}^{\star}\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}+\frac{1}{2C_{0}}\sigma_{r}^{\star}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}},$
@@
$\left\|T_{4}\right\|_{2}\leq$ $\frac{4+3C_{10}}{C_{0}}\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}+\frac{4+3C_{10}}{C_{0}}\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}+\frac{1}{C_{0}}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}.$
@@
$\left(1-\frac{4+3C_{10}}{C_{0}}\right)\left\|\Delta^{t+1,\infty}\right\|_{2,\infty}\leq\frac{20+3C_{10}}{C_{0}}\left(\frac{1}{\mu r}\sqrt{\frac{\mu r}{n}}\right)\gamma^{t+1}+\frac{1}{C_{0}}\left\|D^{t+1,\infty}\right\|_{\mathrm{F}}.$ (17)
@@
$\sum_{k=1}^{n}w_{ik}\left(V^{t+1,m}_{kj}\right)^{2}=$ $\left\|\mathcal{P}_{\Omega_{S^{\star}}^{(-m)}}\left(e_{i}e_{j}^{T}\cdot\left(V^{t+1,m}\right)^{T}\right)\right\|_{\mathrm{F}}^{2}$
$\leq$ $\left\|\mathcal{P}_{\Omega_{S^{\star}}}\left(e_{i}e_{j}^{T}\cdot\left(V^{t+1,m}\right)^{T}\right)\right\|_{\mathrm{F}}^{2}\leq 2\alpha pn\cdot\left\|V^{t+1,m}\right\|_{2,\infty}^{2},$
where the last inequality is due to Lemma 13. Therefore,
@@
$\leq$ $\sum_{i=1}^{n}\left[\sum_{k=1}^{n}w_{ik}\left(S^{t,m}-S^{t}\right)_{ik}^{2}\right]\cdot\left(2\alpha pnr\left\|V^{t+1,m}\right\|_{2,\infty}^{2}\right)$
$=$ $\left\|\mathcal{P}_{\Omega_{S^{\star}}}^{(-m)}\left(S^{t,m}-S^{t}\right)\right\|_{\mathrm{F}}^{2}\cdot\left(2\alpha pnr\left\|V^{t+1,m}\right\|_{2,\infty}^{2}\right).$